# MAGPIE: Multilevel–Adaptive–Guided solver for ptychographic phase retrieval

### v4

@@ Abstract
Ptychography is a state-of-the-art coherent diffractive imaging technique that has attracted significant attention for its ability to reconstruct high-resolution, complex-valued images from a series of diffraction patterns. The technique finds applications in materials science [1, 2], biological imaging [3, 4], integrated circuit imaging [5, 6], and X-ray crystallography [7] (see [8] for a comprehensive overview). In a typical ptychographic experiment, a coherent beam is scanned over overlapping regions of an object while recording the resulting diffraction patterns. Since only the intensity of the diffracted waves is measured—and phase information is lost—the reconstruction task becomes a phase retrieval problem, requiring the recovery of both the amplitude and the phase of the object.
@@
Mathematically, ptychographic phase retrieval is formulated as an inverse problem that is inherently nonconvex and ill-posed. To address these challenges, numerous algorithms have been proposed, including the well-known Ptychographical Iterative Engine (PIE) family [9, 10, 11]. This family of algorithms casts the problem in an iterative fashion and stabilizes the reconstruction through adding a quadratic regularization term.
@@
Another feature of the ptychographic phase retrieval problem is its inherent multiscale structure. Coarse and fine features coexist in the to-be-reconstructed object, requiring different levels of numerical accuracy. This naturally calls for a multigrid solver. Multigrid methods have a long history in the numerical solution of partial differential equations, exploiting the inherent multi-scale structure of the underlying problem by solving it on a hierarchy of discretizations [12]. Coarse-grid approximations capture the global structure of the solution, while fine-grid corrections resolve detailed features. In the context of optimization, a multilevel [^1] Throughout this work, the terms “grid” and “level” are used interchangeably. optimization algorithm (termed MG/OPT) [13] has demonstrated significant improvements in convergence performance, while maintaining flexibility over a wide range of applications [14, 15, 16]. Specific to inverse problems, this hierarchical structure also helps mitigate the ill-posedness inherent in many inverse problems [17, 18, 19]. Recent work has adapted MG/OPT to the ptychographic context [20], resulting in reduced computational costs and enhanced convergence performance.
@@
In this paper, we propose MAGPIE (Multilevel–Adaptive–Guided Ptychographic Iterative Engine), a solver that integrates the advantages of both PIE and MG/OPT. As done in PIE, we recast the problem into an iterative minimization problem with each step solving a quadratic surrogate that majorizes the original objective. The majorization effect brings us a favorable convergence rate. Meanwhile, we also deploy a multigrid-based optimization strategy and reconstruct the coarse and fine features of the object level by level. Unlike other multigrid-based optimization methods whose coarse-level models are generically incompatible with those on the fine level, our proposed surrogate naturally builds compatibility across scales, thereby eliminating the correction term typically required by MG/OPT. This compatibility also ensures automatic hyperparameter selection: once parameters are tuned on the finest grid, those for coarser levels are derived automatically, simplifying tuning in the implementation and improving usability. The overall method is run in a stochastic fashion, and the final algorithm delivers nice reconstruction with substantially reduced cost of full-gradient evaluations and accelerated computational performance.
@@ 1 Preliminaries
In 2D ptychographic phase retrieval, the goal is to reconstruct the object of interest $\bm{z}^{\ast}\in\mathbb{C}^{n^{2}}$ (in vectorized form) from its observed diffraction patterns. The data $\bm{d}_{k}\in\mathbb{R}^{m^{2}}$ are collected as phaseless intensity measurements from overlapping scanning regions. Specifically,
@@
where $N$ is the total number of scans, $P_{k}\in\mathbb{R}^{m^{2}\times n^{2}}$ projects the image onto the $k$-th scanning region, $\bm{Q}\in\mathbb{C}^{m^{2}}$ represents the probe, $\mathcal{F}\in\mathbb{C}^{m^{2}\times m^{2}}$ denotes the two-dimensional discrete Fourier transform operator, $\epsilon_{k}\in\mathbb{R}^{m^{2}}$ corresponds to the noise associated with the $k$-th measurement, $\odot$ denotes the Hadamard (element-wise) product, and $|\cdot|^{2}$ represents the element-wise intensity. The diagonal elements of $P_{k}$ are $1$ for the columns corresponding to the illuminated pixels in the $k$-th scanning region. We also refer to $\bm{Q}\odot P_{k}\bm{z}^{\ast}$ as the $k$-th exit wave. The overlap ratio refers to the fraction of the probe area that overlaps between adjacent scan positions. For example, a 50% overlap ratio means that the probe is shifted by half of its size in each direction when moving between neighboring positions. Higher overlap ratios increase data redundancy and generally improve the stability and convergence of reconstruction algorithms, at the cost of increased measurement acquisition [8]. Figure 1 illustrates the experimental setup and data acquisition process for ptychography.
@@
where $\sqrt{\cdot}$ denotes the element-wise square root. Following the terminology in [11] and invoking Parseval’s theorem, the problem can equivalently be written as:
@@
$=\sum_{k=1}^{N}\Phi_{k}(\bm{z}_{k})=\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\bm{z}_{k}-\mathcal{R}_{k}(\bm{z}_{k})\right\|_{2}^{2}\,,$
where $\bm{z}_{k}$ and the operator $\mathcal{R}_{k}$ are defined below, and the symbol $\odot$ is omitted for notational convenience.
@@
Since the objective function is real-valued but defined on a complex variable, it is not complex-differentiable in the classical sense [21]. A common remedy is to employ the $\mathbb{C}\mathbb{R}$-calculus [22]. To that end, we represent vectors in the complex field, $\bm{z}_{k}\in\mathbb{C}^{m^{2}}$, in the Euclidean space $\mathbb{R}^{2m^{2}}$ as follows:
@@
$=\overline{\bm{Q}}\left(\bm{Q}\bm{z}_{k}-\mathcal{R}_{k}(\bm{z}_{k})\right)\,,$
(See Appendix A.1 for the derivation). Consequently, the full gradient of  Eqn. (1) is
$\nabla_{\bm{z}}\Phi(\bm{z})$ $=\sum_{k=1}^{N}P_{k}^{\top}\nabla_{\bm{z}_{k}}\Phi_{k}(P_{k}\bm{z})$
$=\sum_{k=1}^{N}P_{k}^{\top}\overline{\bm{Q}}\left(\bm{Q}P_{k}\bm{z}-\mathcal{R}_{k}(P_{k}\bm{z})\right)\,.$
@@
The PIE family of algorithms [9, 10, 11] is widely employed to minimize the misfit  Eqn. (1). At each iteration, the scanning regions are randomly shuffled and then updated one by one using a regularized quadratic surrogate model until all regions have been updated. While one scanning region is selected and updated, all entries outside that region remain unchanged. Let $\bm{z}_{k}^{(j)}$ denote the estimate for the $k$-th region at the $j$-th iteration [^3] In each iteration, we shuffle the indices of the scanning regions $\{k\}_{k=1}^{N}$ and update them sequentially.. The local update is then performed by
$\bm{z}_{k}^{(j)}\leftarrow\bm{z}_{k}^{+}=\bm{z}_{k}^{(j)}+\frac{\overline{\bm{Q}}}{\bm{u}+|\bm{Q}|^{2}}\left(\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)-\bm{Q}\bm{z}_{k}^{(j)}\right)\,.$ (3)
Specifically, we define its corresponding regularized quadratic surrogate function $\Phi^{\texttt{PIE}}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)$ as
$\Phi^{\texttt{PIE}}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)=\frac{1}{2}\left\|\bm{Q}\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right\|_{2}^{2}+\frac{1}{2}\bm{u}\cdot\left|\bm{z}_{k}-\bm{z}_{k}^{(j)}\right|^{2}\,,$
@@
where $\bm{u}\in\mathbb{R}^{m^{2}}$ is an a priori chosen regularization coefficient and is applied element-wise to $\left|\bm{z}_{k}-\bm{z}_{k}^{(j)}\right|^{2}$. Then, $\bm{z}^{(j+1)}$ is obtained as the final output of  Eqn. (3) after sweeping through all scanning regions. Notably, for a given $\bm{z}_{k}^{(j)}$, this objective function is quadratic in $\bm{z}_{k}$.
@@
Correspondingly, the complex gradient of $\Phi^{\texttt{PIE}}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)$ with respect to $\bm{z}_{k}$ is:
$\nabla_{\bm{z}_{k}}\Phi^{\texttt{PIE}}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)=\overline{\bm{Q}}\left(\bm{Q}\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)+\bm{u}\left(\bm{z}_{k}-\bm{z}_{k}^{(j)}\right)\,,$ (4)
(See Appendix A.2 for the derivation). Therefore, we set the gradient to zero and solve for the explicit solution $\bm{z}_{k}^{+}$:
$\bm{u}\left(\bm{z}_{k}^{+}-\bm{z}_{k}^{(j)}\right)$ $=\overline{\bm{Q}}\left(\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)-\bm{Q}\bm{z}_{k}^{+}\right)\,,$
$=\overline{\bm{Q}}\left(\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)-\bm{Q}\left(\bm{z}_{k}^{+}-\bm{z}_{k}^{(j)}+\bm{z}_{k}^{(j)}\right)\right)\,,$
implying ${\text{implying that}}\quad\left(\bm{u}+|\bm{Q}|^{2}\right)\left(\bm{z}_{k}^{+}-\bm{z}_{k}^{(j)}\right)$ $=\overline{\bm{Q}}\left(\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)-\bm{Q}\bm{z}_{k}^{(j)}\right)\,,$
recovering the local update rule in Eqn. (3).
@@
The choice of $\bm{u}$ significantly influences the reconstruction accuracy. Within the PIE family of algorithms, experiments in [11] reported that the best performance is achieved by setting:
$\bm{u}=\bm{u}^{\texttt{rPIE}}=\alpha\left(\|\bm{Q}\|^{2}_{\infty}-|\bm{Q}|^{2}\right)\,,$
@@
where $\alpha>0$, and $\|\bm{Q}\|^{2}_{\infty}$ is the maximum intensity among the probe’s entries [^4] In this work, $\bm{Q}$ is flattened into a vector in $\mathbb{C}^{m^{2}}$, so its infinity norm $\|\bm{Q}\|_{\infty}$ equals the maximum magnitude among its entries.. This algorithm is termed regularized Ptychographical Iterative Engine (rPIE). The rationale behind this choice is that in regions of the object with low probe illumination ($|\bm{Q}|_{r}\ll 1$), the revised exit wave is more susceptible to noise, necessitating a larger penalty to stabilize the update. Conversely, where the probe strongly illuminates the object, a smaller penalty is applied, reflecting our increased confidence in those measurements. This adaptive weighting not only enhances the robustness of rPIE but also effectively mitigates noise-induced instability. The pseudocode for rPIE is presented in Algorithm 1. However, despite these advantages, rPIE currently lacks a rigorous theoretical foundation in terms of convergence. We reexamine the algorithm from a surrogate-minimization perspective in Section 2.
@@
3:     for $k$ randomly drawn from $1$ to $N$ without replacement do
4:        $\bm{z}_{k}^{(j)}\leftarrow\bm{z}_{k}^{(j)}+\frac{\overline{\bm{Q}}}{\bm{u}+|\bm{Q}|^{2}}\left(\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)-\bm{Q}\bm{z}_{k}^{(j)}\right)$
@@ 1.2 Multilevel optimization algorithm (MG/OPT)
Ptychographic phase retrieval contains an intrinsic resolution hierarchy, making it an ideal candidate for multigrid acceleration. MG/OPT is a multigrid solver for optimization. Given an objective function $f_{h}$ defined on a fine grid, our goal is to solve $\min_{\bm{z}}f_{h}(\bm{z})$. MG/OPT accelerates this process by, when appropriate, lifting the problem to a coarser level where a simpler optimization problem can be solved more efficiently. By communicating information between fine and coarse grids and offloading expensive fine-grid computations onto the cheaper coarse level, the method speeds up overall convergence. It has already demonstrated faster convergence for the ptychographic phase-retrieval problem [20]. Here, we present the two-level MG/OPT algorithm, which employs one fine grid and one coarse grid; the multilevel extension follows naturally. We introduce the following notation:
@@
  • $\bm{I}_{h}^{H}$: Denotes the restriction operator (binning) that downsamples variables from the fine grid to the coarse grid via average pooling. Specifically,
@@
where the correction term $-\bar{\bm{v}}_{H}^{T}\bm{z}_{H}$ enforces matching of first-order conditions across levels:
@@
The rigorous analysis that validates MG/OPT is given in [13]. Under mild assumptions, the search direction $\bm{e}_{h}$ is a descent direction of $f_{h}$ at $\bar{\bm{z}}_{h}$; that is,

### v6

@@ 1 Introduction
Ptychography is a state-of-the-art coherent diffractive imaging technique that has attracted significant attention for its ability to reconstruct high-resolution, complex-valued images from a series of diffraction patterns. The technique finds applications in materials science [crystallography, materials_science_2], biological imaging [biology_1, biology_2], and integrated circuit imaging [integrated_circuit, integrated_circuit_1](see [Rodenburg2019] for a comprehensive overview). In a typical ptychographic experiment, a coherent beam is scanned over overlapping regions of an object while recording the resulting diffraction patterns. Since only the intensity of the diffracted waves is measured—and phase information is lost—the reconstruction task becomes a phase retrieval problem, requiring the recovery of both the amplitude and the phase of the object.
@@
Mathematically, ptychographic phase retrieval is formulated as an inverse problem that is inherently nonconvex and ill-posed. To address these challenges, numerous algorithms have been proposed, including the well-known Ptychographical Iterative Engine (PIE) family [PIE, ePIE, rPIE]. This family of algorithms casts the problem in an iterative fashion and stabilizes the reconstruction through adding a quadratic regularization term.
@@
Another feature of the ptychographic phase retrieval problem is its inherent multiscale structure. Coarse and fine features coexist in the to-be-reconstructed object, requiring different levels of numerical accuracy. This naturally calls for a multigrid solver. Multigrid methods have a long history in the numerical solution of partial differential equations, exploiting the inherent multi-scale structure of the underlying problem by solving it on a hierarchy of discretizations [multigrid_book]. Coarse-grid approximations capture the global structure of the solution, while fine-grid corrections resolve detailed features. In the context of optimization, a multilevel [^1] Throughout this work, the terms “grid” and “level” are used interchangeably. optimization algorithm (termed MG/OPT) [MGOPT] has demonstrated significant improvements in convergence performance, while maintaining flexibility over a wide range of applications [Nash01012000, mgopt_1, mgopt_2]. Specific to inverse problems, this hierarchical structure also helps mitigate the ill-posedness inherent in many inverse problems [mginverse1, mginverse2, mginverse3]. Recent work has adapted MG/OPT to the ptychographic context [MGOPT_ptycho], resulting in reduced computational costs and enhanced convergence performance.
@@
In this paper, we propose MAGPIE (Multilevel–Adaptive–Guided Ptychographic Iterative Engine), a solver that integrates the advantages of both PIE and MG/OPT. As done in PIE, we recast the problem into an iterative minimization problem with each step solving a quadratic surrogate that majorizes the original objective. The majorization effect brings us a favorable convergence rate. Meanwhile, we also deploy a multigrid-based optimization strategy and reconstruct the coarse and fine features of the object level by level. Unlike other multigrid-based optimization methods whose coarse-level models are generically incompatible with those on the fine level, our proposed surrogate naturally builds compatibility across scales. This compatibility also ensures automatic hyperparameter selection: once parameters are tuned on the finest grid, those for coarser levels are derived automatically, simplifying tuning in the implementation and improving usability. The overall method is run in a stochastic fashion, and the final algorithm delivers nice reconstruction with substantially reduced cost of full-gradient evaluations and accelerated computational performance.
@@ 2 Preliminaries
In 2D ptychographic phase retrieval, the goal is to reconstruct the object $\bm{z}^{\ast}\in\mathbb{C}^{n^{2}}$ (in vectorized form) from measured diffraction intensities. The data $\bm{d}_{k}\in\mathbb{R}^{m^{2}}$ are phaseless intensity measurements collected from overlapping probe positions:
@@
where $N$ is the number of scans, $P_{k}\in\mathbb{R}^{m^{2}\times n^{2}}$ extracts the $k$-th scanning region, $\bm{Q}\in\mathbb{C}^{m^{2}}$ is the probe, $\mathcal{F}$ is the two-dimensional discrete Fourier transform, $\epsilon_{k}$ models noise, $\odot$ denotes element-wise multiplication, and $|\cdot|^{2}$ applies element-wise to yield intensities. We refer to $\bm{Q}\odot P_{k}\bm{z}^{\ast}$ as the $k$-th exit wave. The overlap ratio is the fraction of the probe footprint shared between adjacent scans (e.g., $50\%$ overlap corresponds to half-probe shifts). Higher overlap ratios increase data redundancy and generally improve reconstruction stability and convergence, at the cost of additional measurements [Rodenburg2019]. Figure 1 illustrates the ptychographic acquisition geometry.
@@
where $\sqrt{\cdot}$ and $|\cdot|$ are applied element-wise. Following the terminology in [rPIE] and invoking Parseval’s theorem, the problem can equivalently be written as:
@@
$=\sum_{k=1}^{N}\Phi_{k}(\bm{z}_{k})=\frac{1}{2}\sum_{k=1}^{N}\left\|\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}(\bm{z}_{k})\right\|_{2}^{2}\,$,
where $\bm{z}_{k}$ and the operator $\mathcal{R}_{k}$ are defined below.
@@
Since the objective function is real-valued but defined on a complex variable, it is not complex-differentiable in the classical sense [lang1985complex]. A common remedy is to employ the $\mathbb{C}\mathbb{R}$-calculus [CRcalculus]. To that end, we represent vectors in the complex field, $\bm{z}_{k}\in\mathbb{C}^{m^{2}}$, in the Euclidean space $\mathbb{R}^{2m^{2}}$ as follows:
@@
$=\overline{\bm{Q}}\odot\left(\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}(\bm{z}_{k})\right)\,$,
(See Supplementary Material, Section SM1.1, for the derivation). Consequently, the full gradient of  Eqn. (1) is
$\nabla_{\bm{z}}\Phi(\bm{z})=\sum_{k=1}^{N}P_{k}^{\top}\nabla_{\bm{z}_{k}}\Phi_{k}(P_{k}\bm{z})=\sum_{k=1}^{N}P_{k}^{\top}\left(\overline{\bm{Q}}\odot\left(\bm{Q}\odot P_{k}\bm{z}-\mathcal{R}_{k}\left(P_{k}\bm{z}\right)\right)\right)\,.$
@@
The PIE family of algorithms [PIE, ePIE, rPIE] is widely employed to minimize the misfit  Eqn. (1). At each iteration, the scanning regions are randomly shuffled and then updated one by one using a regularized quadratic surrogate model until all regions have been updated. While one scanning region is selected and updated, all entries outside that region remain unchanged. Let $\bm{z}_{k}^{(j)}$ denote the estimate for the $k$-th region at the $j$-th iteration [^3] In each iteration, we shuffle the indices of the scanning regions $\{k\}_{k=1}^{N}$ and update them sequentially.. The local update is then performed by
$\bm{z}_{k}^{(j)}\leftarrow\bm{z}_{k}^{+}=\bm{z}_{k}^{(j)}+\overline{\bm{Q}}\oslash\left(\bm{u}+|\bm{Q}|^{2}\right)\odot\left(\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)-\bm{Q}\odot\bm{z}_{k}^{(j)}\right)\,$, (3)
[…]
$\Phi^{\texttt{PIE}}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)=\frac{1}{2}\left\|\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right\|_{2}^{2}+\frac{1}{2}\bm{u}^{\top}\left|\bm{z}_{k}-\bm{z}_{k}^{(j)}\right|^{2}\,$,
@@
where $\bm{u}\in\mathbb{R}^{m^{2}}$ is an a priori chosen regularization coefficient. Then, $\bm{z}^{(j+1)}$ is obtained as the final output of  Eqn. (3) after sweeping through all scanning regions. Notably, for a given $\bm{z}_{k}^{(j)}$, this objective function is quadratic in $\bm{z}_{k}$.
@@
Correspondingly, the complex gradient of $\Phi^{\texttt{PIE}}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)$ with respect to $\bm{z}_{k}$ is given by
$\nabla_{\bm{z}_{k}}\Phi^{\texttt{PIE}}_{k}\left(\bm{z}_{k};\bm{z}_{k}^{(j)}\right)=\overline{\bm{Q}}\odot\left(\bm{Q}\odot\bm{z}_{k}-\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)\right)+\bm{u}\odot\left(\bm{z}_{k}-\bm{z}_{k}^{(j)}\right)\,.$ (4)
(see Supplementary Material, Section SM1.2, for the derivation). Therefore, setting the gradient to zero and solving for the explicit solution $\bm{z}_{k}^{+}$ recovers the local update rule in Eqn. (3).
@@
The choice of $\bm{u}$ strongly influences the reconstruction accuracy. Within the PIE family of algorithms, experiments in [rPIE] indicate that the best performance is obtained by choosing
$\bm{u}=\bm{u}^{\texttt{rPIE}}:=\alpha\bigl(\|\bm{Q}\|^{2}_{\infty}\,\bm{J}-|\bm{Q}|^{2}\bigr),$
@@
where $\alpha>0$, $\bm{J}\in\mathbb{R}^{m^{2}}$ denotes the vector with all entries equal to one, and $|\cdot|^{2}$ is applied element-wise. [^4] In this work, $\bm{Q}$ is flattened into a vector in $\mathbb{C}^{m^{2}}$, so its infinity norm $\|\bm{Q}\|_{\infty}$ equals the maximum magnitude among its entries. Equivalently, $\left[\bm{u}^{\texttt{rPIE}}\right]_{r}=\alpha\bigl(\|\bm{Q}\|_{\infty}^{2}-|\bm{Q}_{r}|^{2}\bigr),\quad\text{for }r=1,\dots,m^{2}.$. This algorithm is termed the regularized Ptychographical Iterative Engine (rPIE). The rationale behind this choice is that at indices with low probe illumination ($|\bm{Q}_{r}|\ll 1$), the revised exit wave is more susceptible to noise, necessitating a larger penalty to stabilize the update. Conversely, where the probe strongly illuminates the object, a smaller penalty is applied, reflecting our increased confidence in those measurements. This adaptive weighting not only enhances the robustness of rPIE but also effectively mitigates noise-induced instability. The pseudocode for rPIE is presented in Algorithm 1. However, despite these advantages, rPIE currently lacks a rigorous theoretical foundation in terms of convergence. We reexamine the algorithm from a surrogate-minimization perspective in Section 3.
@@
3:  for $k$ randomly drawn from $1$ to NN without replacement do
4:   $\bm{z}_{k}^{(j)}\leftarrow\bm{z}_{k}^{(j)}+\overline{\bm{Q}}\oslash\left(\bm{u}+|\bm{Q}|^{2}\right)\odot\left(\mathcal{R}_{k}\left(\bm{z}_{k}^{(j)}\right)-\bm{Q}\odot\bm{z}_{k}^{(j)}\right)$
@@ 2.2 Multilevel optimization algorithm (MG/OPT)
Ptychographic phase retrieval contains an intrinsic resolution hierarchy, making it an ideal candidate for multigrid acceleration. MG/OPT is a multigrid solver for optimization. Given an objective function $f_{h}$ defined on a fine grid, our goal is to solve $\min_{\bm{z}}f_{h}(\bm{z})$. MG/OPT accelerates this process by, when appropriate, lifting the problem to a coarser level where a simpler optimization problem can be solved more efficiently. By communicating information between fine and coarse grids and offloading expensive fine-grid computations onto the cheaper coarse level, the method speeds up overall convergence. It has already demonstrated faster convergence for the ptychographic phase-retrieval problem [MGOPT_ptycho]. Here, we present the two-level MG/OPT algorithm, which employs one fine grid and one coarse grid; the multilevel extension follows naturally. We introduce the following notation:
@@
  • $\bm{I}_{h}^{H}$: The restriction operator (binning) that downsamples variables from the fine grid to the coarse grid via average pooling. Specifically, $\bm{I}_{h}^{H}\in\mathbb{R}^{n_{H}^{2}\times n_{h}^{2}}$}, and is given by
@@
where the correction term $\bar{\bm{v}}_{H}^{T}\bm{z}_{H}$ enforces matching of first-order conditions across levels:
@@
The rigorous analysis that validates MG/OPT is given in [MGOPT]. Under mild assumptions, the search direction $\bm{e}_{h}$ is a descent direction of $f_{h}$ at $\bar{\bm{z}}_{h}$; that is,